Complete pats(a) hrough (e) below (a) 0btain a point estimate for the populiation proportion of aduts in the country who believe thuat televisions are a oury hey could do ithout Round to three decimal places as needed) which The sample Round to theee decmal places an (c3 Construct and iterpret a99% cortdence rterval a simple sandom sample, the valof 10, and the less than or equal to 5% of lhe the u bon proporton of aub in the courey uho beleve that televsons areakpurythey colddoitot Sted comedchoce below and nany swer boes ฬาท your doce (Type integers or decimals rounded to three decimal places as needed Use ascending order OA. There a[196 chance the proporton of adits in th outry who belev ethat toe so eauury they coulddowho tes between and ов. we are %contde the proporton of ad nne county who beteve that hie tions aro auuryhey coud de wito t beteen!]ana (Type an integer or a deomal Do not round ) tej Use the results of part (c) to construct a 95% conidence interval for the popuiation proportion of aduts in the country who beleve that television are a necessity The 95% cortonce rtervaisOD Round to hree decimal places as needd)
Answer #1

a)
pcap = 527/1013 = 0.5202

c)
for 95% CI, z-value = 1.96

ME = 1.96*sqrt(0.5202*(1-0.5202)/1013))
= 0.0308

CI = (0.5202 - /+ 0.0308)
= (0.4894, 0.5510)

option B

d)
No supermajority do not think it is a luxury because the CI lies below 0.6

e)
CI = (1- 0.5510, 1- 0.4894)
= (0.449 , 0.5106 )